# Intrinsic training dynamics of deep neural networks

### v1

@@ Conclusion
Since $\phi_{\mathtt{ReLU}}$ is polynomial, we can apply Proposition˜F.1 to obtain an open dense set $\Theta$ on which the dimension of the trace of $\mathbb{W}_{\phi}$ remains constant. By Theorem˜3.8, $\phi_{\mathtt{ReLU}}$ satisfies the Frobenius property. By Proposition˜3.7 every $\theta_{0}\in\Theta$ admits a neighborhood $U$ on which it satisfies the intrinsic recoverability property with respect ot $\phi_{\mathtt{ReLU}}$. By Lemma˜3.2 such a parameter $\theta_{0}$ also satisfies the intrinsic metric property on $U$ with respect to $\phi_{\mathtt{ReLU}}$. ∎
Appendix G Proof of Proposition˜3.10
@@
Let $Y\in\mathbb{R}_{>0}^{n\times m}$. Then there exists a unique pair $(\boldsymbol{\alpha},\boldsymbol{\beta})=:\Gamma(Y)$ of vectors $\boldsymbol{\alpha}\in\mathbb{R}_{>0}^{n}$, $\boldsymbol{\beta}\in\mathbb{R}_{>0}^{m}$ such that
$\boldsymbol{\alpha}^{2}=Y\mathrm{diag}(\boldsymbol{\beta})^{-1}\mathbf{1}_{m},\quad\text{and}\quad{\boldsymbol{\beta}}^{2}=Y^{\top}\mathrm{diag}(\boldsymbol{\alpha})^{-1}\mathbf{1}_{n}.$
@@
$S(\boldsymbol{\beta}):=\sqrt{Y\mathrm{diag}(\boldsymbol{\beta})^{-1}\mathbf{1}_{m}},\quad T(\boldsymbol{\alpha}):=\sqrt{Y^{\top}\mathrm{diag}(\boldsymbol{\alpha})^{-1}\mathbf{1}_{n}}.$
Let $D(a,a^{\prime}):=\|\log(a/a^{\prime})\|_{\infty}$ denote the Thompson metric on $(\mathbb{R}_{+}^{*})^{d}$, where $\mathbb{R}_{+}^{*}$ is the set of positive real numbers. It is known that $\left((\mathbb{R}_{+}^{*})^{d},D\right)$ is a complete metric space. The linear operator $Y$ is 1-Lipschitz with respect to $D$, according to the Birkhoff contraction theorem. Moreover, the square root function is $\frac{1}{2}$-Lipschitz in this metric. Hence, the composition $S\circ T$ is $\frac{1}{4}$-contracting. By the Banach fixed-point theorem, there exists a unique fixed point of $S\circ T$, which implies the existence and uniqueness of the pair $(\boldsymbol{\alpha},\boldsymbol{\beta})$ solving the original equations. ∎
See 3.10
@@
Given the general definition of $\phi_{\mathtt{ReLU}}$ (see e.g. Neyshabur et al. (2015); Stock & Gribonval (2022); Gonon et al. (2023)), we study the factorization map
@@
Let $f:\mathbb{R}^{n\times m}\to\mathbb{R}$ and define the loss $\ell(u,V,w)=f\bigl{(}\phi(u,V,w)\bigr{)}$. Writing $Z=\phi(u,V,w)$ and its gradient $G=\nabla f(Z)$, the gradient-flow ODE (1) $\dot{u}=-\partial_{u}\ell,\;\dot{V}=-\partial_{V}\ell,\;\dot{w}=-\partial_{w}\ell$ is:
@@
Moreover by Corollary˜3.9 there exists conservation laws $\mathbf{h}$ and a function $\Gamma$ such that $\theta=(u,V,w)=\Gamma(\phi(\theta),\mathbf{h}(\theta))=\Gamma(Z,\mathbf{h}(\theta))$ so that $\boldsymbol{\alpha}:=u^{2}$ and $\boldsymbol{\beta}:=w^{2}$ (entrywise multiplication) can both be expressed as functions $\boldsymbol{\alpha}(Z,\mathbf{h}(\theta))$ and $\boldsymbol{\beta}(Z,\mathbf{h}(\theta))$. Below we explicit such conservation laws and characterize properties of $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}$.
Step 3: Conserved quantities and elimination of $\boldsymbol{\alpha},\boldsymbol{\beta}$.
The flow (1) preserves the following $n+m$ conservation laws:
@@
for given constants $\boldsymbol{\lambda}\in\mathbb{R}^{n}$ and $\boldsymbol{\mu}\in\mathbb{R}^{m}$ determined by $\theta_{0}$. Since $V_{ij}=Z_{ij}/(u_{i}w_{j})$, then $(u^{2},w^{2})>0$ is a solution of the coupled system
@@
In vector-matrix form (with entrywise squaring):
$\boldsymbol{\alpha}=u^{2},\quad\boldsymbol{\beta}=w^{2},$
$\boldsymbol{\alpha}^{2}-|Z|^{2}\,\mathrm{diag}(\boldsymbol{\beta})^{-1}\mathbf{1}_{m}-\boldsymbol{\lambda}\odot\boldsymbol{\alpha}=0,\quad\boldsymbol{\beta}^{2}-(|Z|^{2})^{\top}\,\mathrm{diag}(\boldsymbol{\alpha})^{-1}\mathbf{1}_{n}-\boldsymbol{\mu}\odot\boldsymbol{\beta}=0,$
where $|Z|^{2}$ is the elementwise square of $Z$ and $\odot$ is the element-wise product.
Special case $\boldsymbol{\lambda}=0,\boldsymbol{\mu}=0$.
@@
$\boldsymbol{\alpha}^{2}=(|Z|^{2})\,\mathrm{diag}(\boldsymbol{\beta})^{-1}\mathbf{1}_{m},\quad\boldsymbol{\beta}^{2}=(|Z|^{2})^{\top}\,\mathrm{diag}(\boldsymbol{\alpha})^{-1}\mathbf{1}_{n}.$ (19)
By Lemma˜G.1 with $Y=|Z|^{2}$ (possible as $Z_{ij}=u_{i}V_{ij}w_{j}\neq 0$ since $\theta\in\Theta$), the exists a unique solution $(\boldsymbol{\alpha},\boldsymbol{\beta})>0$ of the system (19).
In the scalar case $(n=m=1)$ with $|Z|^{2}=z^{2}$ a scalar, the solution is $\boldsymbol{\alpha}=\boldsymbol{\beta}=(|Z|^{2})^{1/3}=|z|^{2/3}.$
Appendix H Proof of Theorem˜4.3.
See 4.3
@@
In the common eigenbasis, (20) becomes for every $i$
@@
By the first step, one already has that on $\Omega$, for any $i>r$: $\sigma_{i}=\mu_{i}=0$ so that $\sigma_{i}=\sigma_{i}^{-}$, and that for any $i\leq r$, $\sigma_{i}>0$ and $\mu_{i}>0$. Thus $\sqrt{\lambda^{2}+4\mu_{i}}>|\lambda|$, the “-” root is negative, while $P=UU^{\top}$ is positive-semidefinite. Therefore $\sigma_{i}=\sigma_{i}^{+}$ for $i\leq r$. Let us define $\Pi_{ZZ^{\top}}\coloneqq W\operatorname{diag}(\underbrace{1,\cdots,1}_{\times r},0,\cdots,0)W^{\top}$ the orthogonal projector on $\mathrm{range}(ZZ^{\top})$. It follows that:
$P=\Pi_{ZZ^{\top}}\times\left[\frac{\lambda}{2}\mathrm{Id}_{n}+\frac{1}{2}\sqrt{\lambda^{2}\mathrm{Id}_{n}+4\,ZZ^{\top}}\right].$ (21)
@@
Because $Q$ and $T:=Z^{\top}Z$ commute, they share an orthonormal eigenbasis in which (22) reduces to
@@
$Q=\Pi_{Z^{\top}Z}\times\left[-\frac{\lambda}{2}I_{m}+\frac{1}{2}\sqrt{\lambda^{2}\mathrm{Id}_{m}+4\,T}\right],$ (23)
@@
Step 5: Uniqueness and conclusion In both cases (21)–(23) are the only solutions consistent with $UU^{\top}\succeq 0$ and $VV^{\top}\succeq 0$ and with $\mathrm{rank}(Z)=r$ on $\Omega$. Finally one has on $\Omega$:
@@
$=-\Pi_{ZZ^{\top}}\times\left[\frac{\lambda}{2}\mathrm{Id}_{n}+\frac{1}{2}\sqrt{\lambda^{2}\mathrm{Id}_{n}+4\,ZZ^{\top}}\right]\nabla f(X)-\nabla f(X)\Pi_{Z^{\top}Z}\times\left[-\frac{\lambda}{2}I_{m}+\frac{1}{2}\sqrt{\lambda^{2}\mathrm{Id}_{m}+4\,Z^{\top}Z}\right],$
@@
Appendix I Proof of Theorem˜4.4
@@
See 4.4
@@
In light of the necessary condition of Theorem˜2.14 we will first characterize $\mathrm{ker}\partial M(\theta)$ for any $\theta=(U,V)$. Then, with $\mathbf{h}(\theta)=U^{\top}U-V^{\top}V$ and $\phi(\theta)=\phi_{\mathtt{Lin}}(\theta)=UV^{\top}$, we will exhibit a subspace $\mathcal{V}$ of $\mathrm{ker}\partial\mathbf{h}(\theta)\cap\mathrm{ker}\partial\phi(\theta)$ such that $\mathcal{V}\subsetneq\mathrm{ker}\partial M(\theta)$. We will then conclude using the needed calculus and  Theorem˜2.14.
@@
By (11) (with $L=2$, $U_{2}=U,U_{1}=V^{\top}$), one can write $M(\theta)\mathrm{vec}(X)=\mathrm{vec}(UU^{T}X+XVV^{\top})$ for any matrix $X\in\mathbb{R}^{n\times m}$. Using the Kronecker product and the fact that $(A\otimes B)\mathrm{vec}(X)=\mathrm{vec}(BXA^{\top})$, this expression can be rewritten as:
$M(\theta)=\mathrm{Id}_{m}\otimes(UU^{T})+(VV^{\top})\otimes\mathrm{Id}_{n}.$ (24)
Thus differentiating (24) yields that for any $(H,K)$ of the same dimensions as $(U,V)$ we have $\partial M(\theta).(H,K)=\mathrm{Id}_{m}\otimes(UH^{T}+HU^{T})+{(VK^{\top}+KV^{\top})}\otimes\mathrm{Id}_{n}$, and thus: $(H,K)\in\mathrm{ker}\partial M(\theta)$ if and only if $\mathrm{Id}_{m}\otimes(UH^{T}+HU^{T})=-{(VK^{\top}+KV^{\top})}\otimes\mathrm{Id}_{n}$. We now show that
@@
The converse inclusion is clear. We now prove the direct inclusion. Let us consider $(H,K)\in\mathrm{ker}\partial{M}(\theta)$, then one has $\mathrm{Id}_{m}\otimes(UH^{T}+HU^{T})=-{(VK^{\top}+KV^{\top})}\otimes\mathrm{Id}_{n}.$ Still using that $(A\otimes B)\mathrm{vec}(X)=\mathrm{vec}(BXA^{\top})$ and denoting $U^{\prime}\coloneqq UH^{T}+HU^{T}$ and $V^{\prime}\coloneqq{(VK^{\top}+KV^{\top})}$, this exactly means that for any matrix $X\in\mathbb{R}^{n\times m}$ one has $U^{\prime}X=-XV^{\prime\top}$. To conclude, we only need to show that this implies the existence of $\mu,{\mu^{\prime}}\in\mathbb{R}$ such that $U^{\prime}=\mu\mathrm{Id}_{n}\,\text{and}\,V^{\prime}=-{\mu^{\prime}}\mathrm{Id}_{m}$, since the equality $U^{\prime}X=-XV^{\prime\top}$ then also implies $\mu=\mu^{\prime}$. This is immediate if $V^{\prime}=0$ since in this case $U^{\prime}$ must also be equal to zero as $U^{\prime}X=0$ for every $X$. Assume now that $V^{\prime}$ is non-zero so there exists a vector $v$ such that $V^{\prime\top}v\neq 0$. Considering any such $v$ and any vector $u$, and setting $X=uv^{\top}$, we have

### v2

@@ F.2.2 Rank properties
Let us consider a parameter $\theta$ such that $\nabla h_{j}(\theta)\neq 0$ for all hidden neurons $j$.
For each edge $i=\mu\to\nu$ one has
[…]
and similarly for each hidden neuron $j^{\prime}$ with a bias
$\frac{\partial h_{j}}{\partial\theta_{j^{\prime}}}(\theta)=\begin{cases}-2\theta_{j^{\prime}}\text{ if }j=j^{\prime}\\ 0\quad\text{otherwise}\\ \end{cases},$
denoting $h(\theta):=(h_{j}(\theta))_{j}\in\mathbb{R}^{m}$ the Jacobian matrix $\partial h(\theta)\in\mathbb{R}^{m\times D}$ has one or two nonzero entries per row: each row corresponds to one edge or one bias, each column to a hidden neuron, and a row has two nonzero entries at the input and output neurons of the corresponding edge (resp. one nonzero entry at the neuron corresponding to the bias).
By standard linear algebra one can show that $\partial h(\theta)$ has a trivial kernel, hence its column are linearly independent. By continuity this remains the case for the $m$ columns of $\partial h(\theta^{\prime})$ in a neighborhood of $\theta$. This is exactly the definition of independent conservation laws.
We now are able to prove Corollary˜F.8.
Proof.
By Lemma˜F.9, there exists (at least) $m$ independent conservation laws around $\mathbf{1}_{D}$. Let us denote them $h_{i}$, $1\leq i\leq m$. By definition of a conservation law (Definition˜2.7) $h$ of $\phi_{\texttt{ReLU}}$ we have that $\nabla h_{i}(\theta)\in\ker(\partial\phi_{\texttt{ReLU}}(\theta))$, $1\leq i\leq m$, and by definition of the independence of conservation laws, the vectors $\nabla h_{i}(\theta)$, $1\leq i\leq m$, are linearly independent. It follows that
$\mathrm{dim}(\ker(\partial\phi_{\texttt{ReLU}}(\mathbf{1}_{D})))\geq m.$
By the rank-nullity theorem this implies
$\mathrm{rank}(\partial\phi_{\texttt{ReLU}}(\mathbf{1}_{D}))\leq D-m,$
and thus, by Proposition˜F.7, one gets
$\mathrm{rank}(\partial\phi_{\texttt{ReLU}}(\mathbf{1}_{D}))=D-m.\qed$
Corollary F.10.
Consider a DAG ReLU architecture with $D$ parameters and $m$ hidden neurons. Assume that the underlying DAG $G$ is prunable at edge $i$. If $\theta$ satisfies $\theta_{j}\neq 0$ for every $j\neq i$, then $\mathrm{rank}(\partial\phi_{\texttt{ReLU}}(\theta))=\mathrm{rank}(\partial\phi_{\texttt{ReLU}}(\mathbf{1}_{D}))=D-m$.
Proof of Corollary˜F.10.
By Lemma˜F.6, the result holds if $\theta_{i}\neq 0$, hence we focus on the case where $\theta_{i}=0$. Denote $\phi=\phi_{\mathtt{ReLU}}$, and $\mathcal{P}_{i}\coloneqq\{p:p\ni i\}$ the set of all paths containing the edge $i$.
[…]
Thus one can write by blocks:
$\partial\phi(\theta)=\begin{pmatrix}\mathbf{0}&\mathbf{u}\\ {D_{\phi^{\prime}(\theta^{\prime})}}(\theta^{\prime})S^{\prime}{D^{\prime}}_{\theta^{\prime}}^{-1}&\mathbf{0},\end{pmatrix}$ (20)
[…]
$\mathrm{rank}(\partial\phi(\theta))=1+\mathrm{rank}(D_{\phi^{\prime}(\theta^{\prime})}S^{\prime}D_{\theta^{\prime}}^{-1})=1+\mathrm{rank}(S^{\prime}).$
Let us now consider the smaller DAG $G^{\prime}_{i}$ obtained by removing edge $i$ from the graph $G$. Since $G$ is prunable at $i$, by Lemma˜F.2 the set $\mathcal{P}_{i}^{c}$ is exactly the set of all paths of $G^{\prime}_{i}$, so that $S^{\prime}=\partial\phi^{\prime}(\mathbf{1}_{D^{\prime}})$ with $\phi^{\prime}$ the path-lifting associated to $G^{\prime}_{i}$, which is a DAG with $D^{\prime}=D-1$ parameters and $m^{\prime}=m$ hidden neurons. By Corollary˜F.8 we obtain:
[…]
2d step: We now show that $\mathrm{rank}(\partial{\phi}(\theta))\leq\mathrm{rank}(\partial{\phi}(\mathbf{1}_{D}))+1$. Indeed
[…]
where in the last inequality we used again Equation˜20, recalling that $\theta_{i}$ is the only coordinate of $\theta$.
[…]
Corollary F.11.
Consider a DAG ReLU architecture with $D$ parameters that is prunable everywhere. For any $\theta$ with at most one zero component, one has $\mathrm{rank}(\partial\phi_{\texttt{ReLU}}(\theta))=\mathrm{rank}(\partial\phi_{\texttt{ReLU}}(\mathbf{1}_{D}))=D-m$, with $m$ the number of hidden neurons.
Proof.
By Corollary˜F.8 we have $\mathrm{rank}(\partial\phi_{\texttt{ReLU}}(\mathbf{1}_{D}))=D-m$ with $m$ the number of hidden neurons and $D$ the number of parameters. Since $\theta$ has at most one zero component and since the underlying graph is prunable at edge $i$ (as it is everywhere by assumption), we conclude using Corollary˜F.10. ∎
In particular, for an MLP architecture, this corollary reads:
Corollary F.12.
Consider an MLP architecture with $D$ parameters where every layer (including input and output ones) contains at least two neurons. For any $\theta$ with at most one zero component, one has $\mathrm{rank}(\partial\phi_{\texttt{ReLU}}(\theta))=\mathrm{rank}(\partial\phi_{\texttt{ReLU}}(\mathbf{1}_{D}))=D-m$, with $m$ the number of hidden neurons.
Proof.
By Corollary˜F.8 we have $\mathrm{rank}(\partial\phi_{\texttt{ReLU}}(\mathbf{1}_{D}))=D-m$ with $m$ the number of hidden neurons and $D$ the number of parameters. Since $\theta$ has at most one zero component, there is an edge $i$ in the graph $G$ underlying the MLP architecture such that $\theta_{j}\neq 0$ for each $j\neq i$. Moreover, since we focus on an MLP, by Corollary˜F.5 and our assumption on the layer widths, the underlying graph is prunable everywhere, hence it is prunable at edge $i$. We conclude using Corollary˜F.10. ∎
F.2.3 Proof of Theorem˜3.2
We can now establish Theorem˜3.2. In fact, we prove a more general version of Theorem˜3.2:
Theorem F.13.
Consider a DAG ReLU architecture that is prunable everywhere, and denote $\Theta\subseteq\mathbb{R}^{D}$ the set of all parameters with at most one zero component. On this set, which is open, dense, and with a single connected component, the parameterization $\phi_{\mathtt{ReLU}}$ of Gonon et al. (2024) satisfies the Frobenius property.
Proof.
It is each to check that $\Theta$ is an dense open set with a single connected component. Moreover for any $\theta\in\Theta$, denoting $\phi=\phi_{\mathtt{ReLU}}$, we have
[…]
To conclude we need to prove that this is indeed an equality. Assume by contradiction that there exists a parameter $\theta_{1}\in\Theta$ such that $\mathrm{dim}\mathrm{Lie}(\mathbb{W}_{\phi})(\theta_{1})>d$. Then, by standard continuity arguments, this will remain the case on an open neighborhood $U$ of $\theta_{1}$, which is absurd by Theorem˜3.1 as the Frobenius propery holds on $(\mathbb{R}\setminus\{0\})^{D}$, a dense open subset of $\Theta$. ∎
F.3 Proof of Corollary˜3.3
We prove a more general version of Corollary˜3.3.
Corollary F.14.
For a general DAG ReLU network (resp. for a DAG ReLU network prunable everywhere), every element $\theta$ of the set $(\mathbb{R}\setminus\{0\})^{D}$ (resp. of the set $\Theta$ of Theorem˜3.2/Theorem˜F.13) admits an open neighborhood $U$ in this set on which the intrinsic recoverability property (and thus the intrinsic dynamic property with respect to $\phi_{\mathtt{ReLU}}$) holds.
Proof.
@@
Let us consider a DAG ReLU network (resp. a DAG ReLU architecture prunable everywhere). We can apply Lemma˜F.6 (resp. Corollary˜F.11) to obtain an open dense set $\Theta$ (resp. with a single connected component) on which the dimension of the trace of $\mathbb{W}_{\phi}$ remains constant. By Theorem˜3.1 (resp. Theorem˜F.13), $\phi_{\mathtt{ReLU}}$ satisfies the Frobenius property on $\Theta$. By Proposition˜2.21 every $\theta_{0}\in\Theta$ admits a neighborhood $U$ on which it satisfies the intrinsic recoverability property with respect to $\phi_{\mathtt{ReLU}}$. By Lemma˜2.16 such a parameter $\theta_{0}$ also satisfies the intrinsic metric property on $U$ with respect to $\phi_{\mathtt{ReLU}}$. ∎
F.4 Proof of Corollary˜3.4.
We prove a more general version of Corollary˜3.4.
Corollary F.15.
Let us consider a DAG ReLU architecture that is prunable everywhere, and with $m$ hidden neurons. Then the $m$ independent conservation laws for $\phi_{\texttt{ReLU}}$ given in Proposition˜2.9 are exhaustive on the set $\Theta$ from Theorem˜F.13: there is no more conservation laws than these ones.
Proof.
On $\Theta$, we know that the dimension of $\mathrm{Lie}(\mathbb{W}_{\phi_{\texttt{ReLU}}})(\theta)$ is constant and equal to $\mathrm{range}(\partial{\phi_{\texttt{ReLU}}}(\mathbf{1}_{D}))=D-m$ by Theorem˜3.2 and Corollary˜F.11. Then by (Marcotte et al., 2023, Theorem 3.3), there exists exactly $m$ independent conservation laws locally on $\Theta$. Thus the $m$ independent conservation laws for $\phi_{\texttt{ReLU}}$ on $\Theta$ given in Lemma˜F.9 are exhaustive. ∎
Appendix G Proof of Proposition˜3.5
@@
Let $Y\in\mathbb{R}_{>0}^{n\times m}$. Then there exists a unique pair $(\bm{\alpha},\bm{\beta})=:\Gamma(Y)$ of vectors $\bm{\alpha}\in\mathbb{R}_{>0}^{n}$, $\bm{\beta}\in\mathbb{R}_{>0}^{m}$ such that
$\bm{\alpha}^{2}=Y\mathrm{diag}(\bm{\beta})^{-1}\mathbf{1}_{m},\quad\text{and}\quad{\bm{\beta}}^{2}=Y^{\top}\mathrm{diag}(\bm{\alpha})^{-1}\mathbf{1}_{n}.$
@@
$S(\bm{\beta}):=\sqrt{Y\mathrm{diag}(\bm{\beta})^{-1}\mathbf{1}_{m}},\quad T(\bm{\alpha}):=\sqrt{Y^{\top}\mathrm{diag}(\bm{\alpha})^{-1}\mathbf{1}_{n}}.$
Let $D_{n}(a,a^{\prime}):=\|\log(a/a^{\prime})\|_{\infty}$ denote the Thompson metric on $(\mathbb{R}_{>0})^{n}$ (and similarly $D_{m}$ on $(\mathbb{R}_{>0})^{m}$), where $\mathbb{R}_{+}^{*}$ is the set of positive real numbers. It is known that $\left((\mathbb{R}_{>0})^{n},D_{n}\right)$ (resp. $\left((\mathbb{R}_{>0})^{m},D_{m}\right)$) is a complete metric space. The linear operator $Y$ is 1-Lipschitz from $\left((\mathbb{R}_{>0})^{m},D_{m}\right)$ to $\left((\mathbb{R}_{>0})^{n},D_{n}\right)$, according to the Birkhoff contraction theorem. Moreover, the square root function is $\frac{1}{2}$-Lipschitz in this metric. Hence, the composition $S\circ T$ is $\frac{1}{4}$-contracting with respect to $D_{n}$. By the Banach fixed-point theorem, there exists a unique fixed point of $S\circ T$, which implies the existence and uniqueness of the pair $(\bm{\alpha},\bm{\beta})$ solving the original equations. ∎
See 3.5
@@
Given the general definition of $\phi_{\mathtt{ReLU}}$ (see e.g. Neyshabur et al. (2015); Stock and Gribonval (2022); Gonon et al. (2024)), we study the factorization map
@@
Let $f:\mathbb{R}^{n\times m}\to\mathbb{R}$ and define the loss $\ell(u,V,w)=f\bigl(\phi(u,V,w)\bigr)$. Writing $Z=\phi(u,V,w)$ and its gradient $G=\nabla f(Z)$, the gradient‑flow ODE equation 1 $\dot{u}=-\partial_{u}\ell,\;\dot{V}=-\partial_{V}\ell,\;\dot{w}=-\partial_{w}\ell$ is:
@@
Moreover by Corollary˜3.3 there exists conservation laws $\mathbf{h}$ and a function $\Gamma$ such that $\theta=(u,V,w)=\Gamma(\phi(\theta),\mathbf{h}(\theta))=\Gamma(Z,\mathbf{h}(\theta))$ so that $\bm{\alpha}:=u^{2}$ and $\bm{\beta}:=w^{2}$ (entrywise multiplication) can both be expressed as functions $\bm{\alpha}(Z,\mathbf{h}(\theta))$ and $\bm{\beta}(Z,\mathbf{h}(\theta))$. Below we explicit such conservation laws and characterize properties of $\bm{\alpha}$ and $\bm{\beta}$.
Step 3: Conserved quantities and elimination of $\bm{\alpha},\bm{\beta}$.
The flow equation 1 preserves the following $n+m$ conservation laws:
@@
for given constants $\bm{\lambda}\in\mathbb{R}^{n}$ and $\bm{\mu}\in\mathbb{R}^{m}$ determined by $\theta_{0}$. Since $V_{ij}=Z_{ij}/(u_{i}w_{j})$, then $(u^{2},w^{2})>0$ is a solution of the coupled system
@@
In vector‑matrix form (with entrywise squaring):
$\bm{\alpha}=u^{2},\quad\bm{\beta}=w^{2},$
$\bm{\alpha}^{2}-|Z|^{2}\,\mathrm{diag}(\bm{\beta})^{-1}\mathbf{1}_{m}-\bm{\lambda}\odot\bm{\alpha}=0,\quad\bm{\beta}^{2}-(|Z|^{2})^{\top}\,\mathrm{diag}(\bm{\alpha})^{-1}\mathbf{1}_{n}-\bm{\mu}\odot\bm{\beta}=0,$
where $|Z|^{2}$ is the elementwise square of $Z$ and $\odot$ is the element‑wise product.
Special case $\bm{\lambda}=0,\bm{\mu}=0$.
@@
$\bm{\alpha}^{2}=(|Z|^{2})\,\mathrm{diag}(\bm{\beta})^{-1}\mathbf{1}_{m},\quad\bm{\beta}^{2}=(|Z|^{2})^{\top}\,\mathrm{diag}(\bm{\alpha})^{-1}\mathbf{1}_{n}.$ (21)
By Lemma˜G.1 with $Y=|Z|^{2}$ (possible as $Z_{ij}=u_{i}V_{ij}w_{j}\neq 0$ since $\theta\in\Theta$), the exists a unique solution $(\bm{\alpha},\bm{\beta})>0$ of the system equation 21.
In the scalar case $(n=m=1)$ with $|Z|^{2}=z^{2}$ a scalar, the solution is $\bm{\alpha}=\bm{\beta}=(|Z|^{2})^{1/3}=|z|^{2/3}.$
Appendix H Proof of Theorem˜3.8.
See 3.8
@@
In the common eigenbasis, equation 22 becomes for every $i$
@@
By the first step, one already has that on $\Omega$, for any $i>r$: $\sigma_{i}=\mu_{i}=0$ so that $\sigma_{i}=\sigma_{i}^{-}$, and that for any $i\leq r$, $\sigma_{i}>0$ and $\mu_{i}>0$. Thus $\sqrt{\lambda^{2}+4\mu_{i}}>|\lambda|$, the “-” root is negative, while $P=UU^{\top}$ is positive‑semidefinite. Therefore $\sigma_{i}=\sigma_{i}^{+}$ for $i\leq r$. Let us define $\Pi_{Z}\coloneqq W\operatorname{diag}(\underbrace{1,\cdots,1}_{\times r},0,\cdots,0)W^{\top}$ the orthogonal projector on $\mathrm{range}(Z)=\mathrm{range}(ZZ^{\top})$. It follows that:
$P=\Pi_{Z}\times\left[\frac{\lambda}{2}\mathrm{Id}_{n}+\frac{1}{2}\sqrt{\lambda^{2}\mathrm{Id}_{n}+4\,ZZ^{\top}}\right].$ (23)
@@
Because $Q$ and $T:=Z^{\top}Z$ commute, they share an orthonormal eigenbasis in which equation 24 reduces to
@@
$Q=\Pi_{Z^{\top}Z}\times\left[-\frac{\lambda}{2}\mathrm{Id}_{m}+\frac{1}{2}\sqrt{\lambda^{2}\mathrm{Id}_{m}+4\,T}\right],$ (25)
@@
Step 5: Uniqueness and conclusion In both cases equation 23–equation 25 are the only solutions consistent with $UU^{\top}\succeq 0$ and $VV^{\top}\succeq 0$ and with $\mathrm{rank}(Z)=r$ on $\Omega$. Finally one has on $\Omega$:
@@
$=-\Pi_{Z}\left[\frac{\lambda}{2}\mathrm{Id}_{n}+\frac{1}{2}\sqrt{\lambda^{2}\mathrm{Id}_{n}+4\,ZZ^{\top}}\right]\nabla f(X)-\nabla f(X)\Pi_{Z^{\top}Z}\left[-\frac{\lambda}{2}\mathrm{Id}_{m}+\frac{1}{2}\sqrt{\lambda^{2}\mathrm{Id}_{m}+4\,Z^{\top}Z}\right],$
@@
Appendix I Proof of Theorem˜3.9
@@
We introduce a slightly stronger version of Definition˜2.6 involving piecewise $\mathcal{C}^{2}$ functions $f$.
Definition I.2 (Strong intrinsic dynamic property).
$\theta_{0}$ verifies the strong intrinsic dynamic property with respect to $\phi$ on some open set $\Omega\ni\theta_{0}$, if there is $K_{\theta_{0}}:\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ such that: if $\theta(\cdot)\in\mathcal{C}^{0}$ satisfies $\theta(0)=\theta_{0}$ and $\dot{\theta}(t)=-\nabla\ell(\theta(t))$ with $\ell=f\circ\phi$, where $f$ is piecewise $\mathcal{C}^{2}$, whenever $\theta(t)\in\Omega$ and $f$ is differentiable at $\phi(\theta(t))$, then $M(\theta(t))=K_{\theta_{0}}(\phi(\theta(t)))$ holds for each $t$ such that $\theta(t)\in\Omega$.
See 3.9
@@
A. First we show that such an initialization does not satisfy the intrinsic metric property. In light of the necessary condition of Theorem˜2.14 we will first characterize $\mathrm{ker}\partial M(\theta)$ for any $\theta=(U,V)$. Then, with $\mathbf{h}(\theta)=U^{\top}U-V^{\top}V$ and $\phi(\theta)=\phi_{\mathtt{Lin}}(\theta)=UV^{\top}$, we will exhibit a subspace $\mathcal{V}$ of $\mathrm{ker}\partial\mathbf{h}(\theta)\cap\mathrm{ker}\partial\phi(\theta)$ such that $\mathcal{V}\not\subset\mathrm{ker}\partial M(\theta)$. We will then conclude using the needed calculus and  Theorem˜2.14.
@@
By equation 11 (with $L=2$, $U_{2}=U,U_{1}=V^{\top}$), one can write $M(\theta)\mathrm{vec}(X)=\mathrm{vec}(UU^{\top}X+XVV^{\top})$ for any matrix $X\in\mathbb{R}^{n\times m}$. Using the Kronecker product and the fact that $(A\otimes B)\mathrm{vec}(X)=\mathrm{vec}(BXA^{\top})$, this expression can be rewritten as:
$M(\theta)=\mathrm{Id}_{m}\otimes(UU^{\top})+(VV^{\top})\otimes\mathrm{Id}_{n}.$ (26)
Thus differentiating equation 26 yields that for any $(H,K)$ of the same dimensions as $(U,V)$ we have $\partial M(\theta).(H,K)=\mathrm{Id}_{m}\otimes(UH^{\top}+HU^{\top})+{(VK^{\top}+KV^{\top})}\otimes\mathrm{Id}_{n}$, and thus: $(H,K)\in\mathrm{ker}\partial M(\theta)$ if and only if $\mathrm{Id}_{m}\otimes(UH^{\top}+HU^{\top})=-{(VK^{\top}+KV^{\top})}\otimes\mathrm{Id}_{n}$. We now show that
@@
The converse inclusion is clear. We now prove the direct inclusion. Let us consider $(H,K)\in\mathrm{ker}\partial{M}(\theta)$, then one has $\mathrm{Id}_{m}\otimes(UH^{\top}+HU^{\top})=-{(VK^{\top}+KV^{\top})}\otimes\mathrm{Id}_{n}.$ Still using that $(A\otimes B)\mathrm{vec}(X)=\mathrm{vec}(BXA^{\top})$ and denoting $U^{\prime}\coloneqq UH^{\top}+HU^{\top}$ and $V^{\prime}\coloneqq{(VK^{\top}+KV^{\top})}$, this exactly means that for any matrix $X\in\mathbb{R}^{n\times m}$ one has $U^{\prime}X=-XV^{\prime\top}$. To conclude, we only need to show that this implies the existence of $\mu,{\mu^{\prime}}\in\mathbb{R}$ such that $U^{\prime}=\mu\mathrm{Id}_{n}\,\text{and}\,V^{\prime}=-{\mu^{\prime}}\mathrm{Id}_{m}$, since the equality $U^{\prime}X=-XV^{\prime\top}$ then also implies $\mu=\mu^{\prime}$. This is immediate if $V^{\prime}=0$ since in this case $U^{\prime}$ must also be equal to zero as $U^{\prime}X=0$ for every $X$. Assume now that $V^{\prime}$ is non-zero so there exists a vector $v$ such that $V^{\prime\top}v\neq 0$. Considering any such $v$ and any vector $u$, and setting $X=uv^{\top}$, we have